Sally has 5 more quarters than dimes. She has half as many nickels as dimes . She has an equal number of loonies as dimes. In total she has $15.00. How many quarters does sally have?
Answer by mananth(16949)

Dimes ------x ------- 10
quarters ---x+5---------25
Nickels-----x/2---------5
Loonies------x---------- 100
...
10x+25(x+5)+5x/2+100x= 1500
10x+25x+125+5x/2+100x=1500
multiply by 2
20x+50x+250+5x+200x=3000
275x=3000-250
275x=2750
x= 10 dimes
Quarters = 10+5 = 15 coins